next year LPA goes to washington dc, there will be 195 students going on the trip. there will be 7 vehicles, some buses, some cars. a car can hold 5 students and a bus can hold 45 students. how many of each type of vehicle went on the trip?
Answer by mananth(16946)


next year LPA goes to washington dc,
let there be x buses and y cars
x+y =7 ( Total vehicles)

there will be 195 students going on the trip,. a car can hold 5 students and a bus can hold 45 students.
45x +5y=195

1.00 x + 1.00 y = 7.00


45.00 x + 5.00 y = 195.00 .............2
Eliminate y
multiply (1)by -5.00
Multiply (2) by 1.00
-5.00 x -5.00 y = -35.00
45.00 x 5.00 y = 195.00
Add the two equations
40.00 x = 160.00
/ 40.00
x = 4.00
plug value of x in (1)
1.00 x + 1.00 y = 7.00
4.00 + 1.00 y = 7.00
1.00 y = 3.00
y = 3.00
Ans x = 4 (buses)
y = 3 (cars)